I need help understanding how to solve by graphing:
3x + 4y = 20
3x - 2y = 8

Answer by mananth(16946)

3x + 4y = 20
4y=-3x+20
y=(-3x/4)+5
x=0, y=5 (0,5)
y=0,x=20/3 (20/3,0)
..
3x - 2y = 8
2y=3x-8
y=(3x/2)-4
x=0, y=-4 (0,-4)
y=0, x=8/3 (8/3,0)
....

The solution is the point of intersection (4,2)